### v1

# SEDM: Scalable Self-Evolving Distributed Memory for Agents

Abstract
Long-term multi-agent systems inevitably generate vast amounts of trajectories and historical interactions, which makes efficient memory management essential for both performance and scalability. Existing methods typically depend on vector retrieval and hierarchical storage, yet they are prone to noise accumulation, uncontrolled memory expansion, and limited generalization across domains. To address these challenges, we present SEDM, Self-Evolving Distributed Memory, a verifiable and adaptive framework that transforms memory from a passive repository into an active, self-optimizing component. SEDM integrates verifiable write admission based on reproducible replay, a self-scheduling memory controller that dynamically ranks and consolidates entries according to empirical utility, and cross-domain knowledge diffusion that abstracts reusable insights to support transfer across heterogeneous tasks. Evaluations on benchmark datasets demonstrate that SEDM improves reasoning accuracy while reducing token overhead compared with strong memory baselines, and further enables knowledge distilled from fact verification to enhance multi-hop reasoning. The results highlight SEDM as a scalable and sustainable memory mechanism for open-ended multi-agent collaboration. The code will be released in the later stage of this project.
1 Introduction
In recent years, the rapid development of large-scale multi-agent systems (MAS) Wooldridge (2009); Lowe et al. (2017); Foerster et al. (2016); Yang et al. (2018a); Buşoniu et al. (2008). has expanded their application in diverse domains, including collaborative reasoning, decision-making, and autonomous planning Shoham et al. (2007). A central challenge in these open-ended and long-term tasks is the ability of agents to manage, interpret, and reuse information accumulated from continuous interaction with both peers and the environment Yang et al. (2018a). Without effective memory mechanisms, the sheer scale of historical interactions can easily overwhelm computational resources and compromise decision quality Foerster et al. (2018).
In open-ended and long-term multi-agent tasks, each agent relies on its past memories, the observed states of other agents, and the current environment to determine subsequent actions or responses Das et al. (2017). During continuous interaction between agents and their environment, the MAS gradually accumulates extensive logs of interactions, invocation trajectories, and high-level policy memories OpenAI (2023). Such overwhelming amounts of information directly impact the efficiency and cost of decision-making, often leading to higher monetary costs and longer contextual requirements for inference Shuster et al. (2020). Therefore, designing an efficient and sustainable memory mechanism has become a critical issue for modern long-term multi-agent systems.
Current methods primarily adopt vector retrieval and hierarchical memory structures to manage storage and retrieval efficiently Johnson et al. (2019b). Vector retrieval Johnson et al. (2019a); Karpukhin et al. (2020); Izacard and Grave (2021); Lewis et al. (2020); Guo et al. (2016) leverages semantic similarity to identify relevant entries, while hierarchical organization arranges information in layered structures according to abstraction levels Rae et al. (2021). These approaches have shown promise in improving retrieval accuracy and managing memory scalability Chen et al. (2016). However, in complex collaborative multi-agent tasks, their effectiveness diminishes, as the underlying assumptions of stability and linear growth do not hold Wang et al. (2020). This gap between theoretical promise and practical performance highlights several critical limitations that hinder their long-term applicability.
One major challenge is the inevitable accumulation of noise, which severely degrades retrieval quality Fan et al. (2018). As the memory size expands without constraint, the system faces exponentially increasing computational costs in both retrieval and context construction Izacard and Grave (2021). This not only reduces overall efficiency but also amplifies the interference caused by redundant information Lewis et al. (2020). In particular, the presence of low-value or semantically irrelevant entries dilutes the contribution of high-quality information in retrieval results, impairing downstream task performance and leading to measurable declines in metrics Chen et al. (2017). In addition, the cumulative noise effect increases response latency and accelerates the nonlinear consumption of computational and storage resources Karpukhin et al. (2020), ultimately threatening both scalability and stability in long-term MAS operations Press et al. (2020).
To overcome these limitations, we introduce Scalable Self-Evolving Distributed Memory (SEDM), a framework that transforms memory from a passive repository into an adaptive, self-optimizing, and verifiable component for multi-agent systems. Unlike conventional designs that treat memory as a static store, SEDM continually refines knowledge to enhance learning and decision-making efficiency in dynamic task environments. It operationalizes memory as an active mechanism by integrating verifiability and continuous self-improvement into the memory lifecycle. At its core, memory items undergo a rigorous admission process based on self-contained execution contexts (SCECs), such as Docker and ReproZip Merkel (2014); Chirigati et al. (2016), which package all necessary information for environment-free replay and offline validation. This mechanism provides empirical evidence for utility at write time, ensuring that only useful, high-quality experiences enter the memory repository. Once admitted, memory items are dynamically managed by a self-scheduling controller and enhanced through cross-domain knowledge diffusion. The controller leverages admission-derived weights, combined with semantic similarity, to schedule retrieval-time usage without costly reranking, while consolidation and progressive evolution continuously refine the repository by promoting stable items, merging redundancies, and pruning harmful ones. Beyond single-task settings, SEDM abstracts reusable insights into general forms, enabling knowledge distilled in one domain to be safely transferred and re-validated in others. Together, these components establish a scalable and auditable memory mechanism that enhances reasoning accuracy, reduces overhead, and supports sustainable long-term multi-agent collaboration.
We evaluate SEDM on two representative benchmarks, FEVER Thorne et al. (2018) for fact verification and HotpotQA Yang et al. (2018b) for multi-hop reasoning, comparing against no-memory and G-Memory baselines Zhang et al. (2025). The results show that SEDM consistently improves task accuracy while significantly reducing token overhead, thereby achieving a better balance between performance and efficiency. Ablation studies confirm that both the verifiable admission mechanism and the self-scheduling controller contribute progressively to this gain, with the latter playing a key role in constraining prompt growth without sacrificing accuracy. Furthermore, cross-domain evaluation demonstrates that memory distilled from one dataset can transfer to another, with factual knowledge from FEVER notably boosting performance on HotpotQA. These findings highlight SEDM as a scalable, adaptive, and generalizable memory framework for long-term multi-agent reasoning.
Our contributions are summarized as follows:
  • We propose Scalable Self-Evolving Distributed Memory (SEDM), a novel framework that transforms memory from a passive repository into an adaptive, verifiable, and continuously improving component, introducing self-contained execution contexts (SCECs) for reproducible admission and utility-based memory weighting.
  • We design a self-scheduling memory controller that selectively manages memory at retrieval time and continuously refines the repository through consolidation, redundancy suppression, and progressive evolution, thereby balancing accuracy and efficiency.
  • We conduct extensive evaluations on FEVER and HotpotQA benchmarks, demonstrating that SEDM consistently improves task accuracy while significantly reducing token overhead.
2 Related Work
Self-Evolving Agents.
Recent efforts in building self-evolving agents have focused on enabling systems to improve their reasoning or behavior over time without explicit retraining. Approaches such as Reflexion Shinn et al. (2023) and Voyager Wang et al. (2023) allow agents to iteratively refine their strategies by leveraging self-reflection and accumulated trajectories. Similarly, MEMIT Meng et al. (2022) demonstrates the feasibility of localized knowledge editing within large language models, suggesting a pathway for agents to evolve by continuously updating their internal representations. These studies highlight the importance of mechanisms that support autonomous adaptation and progressive self-improvement in dynamic environments.
Agent Memory.
In parallel, research on agent memory has investigated how to store, retrieve, and utilize knowledge efficiently across long-horizon interactions. Episodic memory systems, such as those proposed by Park et al. Park et al. (2023), emulate human-like memory consolidation to support consistent long-term behavior in simulated social environments. Memory-augmented neural networks Graves et al. (2016) and differentiable neural dictionaries Kaiser et al. (2017) further demonstrate how structured memory access can enhance reasoning and generalization. More recently, retrieval-augmented generation frameworks tailored for interactive agents Mialon et al. (2023) have shown that dynamically grounding responses in curated external memories improves both interpretability and task success. Together, these works underscore the need for memory systems that are not only scalable but also adaptive to the agent’s evolving operational context.
3 Methodology
3.1 System Overview
Figure 1: Illustration of different memory strategies. No Memory: the agent interacts with the environment without retaining past information. Fixed Memory: the agent retrieves from a static memory pool, which may grow excessively. SEDM: introduces verifiable write admission, parallel simulation, and adaptive scheduling to build high-quality, self-evolving memory that supports efficient and transferable knowledge use.
Figure 1 illustrates the differences between no memory, fixed memory, and our proposed SEDM framework, highlighting how SEDM achieves verifiable admission, adaptive scheduling, and sustainable knowledge evolution.
Figure 2 gives an end-to-end view of SEDM. The system introduces verifiability and self-improvement into the memory life cycle and consists of three tightly integrated modules. (i) SCEC-based Verifiable Write Admission packages each run into a Self-Contained Execution Context (SCEC) and performs environment-free A/B replay to estimate the marginal utility of a candidate memory item; only items with positive evidence are admitted and assigned an initial weight. (ii) Self-Scheduling in the Memory Controller uses admission-derived weights together with semantic similarity to score candidates at retrieval time, while also maintaining the repository by updating weights from observed outcomes, merging near duplicates, and pruning harmful entries. (iii) Cross-Domain Knowledge Diffusion abstracts admitted items into conservative general forms and re-validates them in other tasks, allowing knowledge to transfer safely across domains.
[…]
3.2 SCEC-based Verifiable Write Admission
We formulate write admission as a verifiable, environment-free procedure that assigns an initial utility weight to each candidate memory item before it enters the repository. The process is based on a Self-Contained Execution Context (SCEC), a minimal and standardized package that enables validation, parallel replay, and offline auditing. By placing admission behind paired A/B evaluations within SCECs, the system produces reproducible evidence for weight initialization while filtering out negative or noisy experiences.
3.2.1 Self-Contained Packaging and Distributed Replay
Each task execution is encapsulated into an SCEC to support reproducible validation and analysis without requiring the original environment. An SCEC includes all necessary inputs, outputs, tool summaries, seeds, and configuration hashes, ensuring (i) self-contained representation, (ii) environment-free replay by summarizing external tool calls, (iii) deterministic reproduction across model versions and seeds, and (iv) minimal sufficiency by storing only essential information.
Treating an SCEC as an independent job enables large-scale distributed A/B replay on arbitrary workers. Only aggregated statistics, along with integrity hashes and version stamps, are uploaded, preserving auditability while controlling cost. This environment-free design eliminates the need to reconstruct complex environments or interact with real agents during validation, thereby allowing memory effectiveness to be tested through parallel replay at scale. As a result, admission decisions can be made rapidly and consistently, significantly reducing computational overhead while ensuring that only high-quality experiences enter the memory repository.
3.2.2 SCEC-grounded A/B Test for Memory Item Initialization
From each SCEC, we extract one candidate memory item $m$, represented as a concise, independently injectable snippet. The extraction process identifies decisive reasoning or corrective steps, performs deduplication and canonicalization, and attaches provenance information.
To evaluate its utility, we conduct a paired A/B test within the same SCEC. The control condition (A) uses the original prompt, while the treatment condition (B) augments the prompt with the candidate memory $m$. This setup isolates the marginal effect of $m$ and provides empirical evidence for its contribution. For a query $q$, the constructed prompts are defined as
$I_{A}=f(q),\qquad I_{B}=f(q;m),$ (1)
where $f(\cdot)$ denotes prompt construction and $I_{B}$ injects $m$ into the SCEC’s dedicated slot together with summarized tool feedback. The model execution inside the SCEC is denoted by $\mathcal{F}$:
$o_{A}=\mathcal{F}(I_{A}),\qquad o_{B}=\mathcal{F}(I_{B}).$ (2)
We then measure the deltas in reward, latency, and token usage:
$\Delta R=R(o_{B})-R(o_{A}),\quad\Delta L=L(o_{B})-L(o_{A}),\quad\Delta T=T(o_{B})-T(o_{A}),$ (3)
where $R(\cdot)$ is the task-specific reward, and $L(\cdot)$ and $T(\cdot)$ denote latency and token overhead, respectively. A composite admission score balances utility and cost:
$S=\Delta R-\lambda_{L}\Delta L-\lambda_{T}\Delta T,$ (4)
with $\lambda_{L},\lambda_{T}\geq 0$ controlling the trade-offs.
The admission decision and initial weight are then defined as
$\mathrm{accept}(m)\Longleftrightarrow S\geq\eta,\qquad w_{0}(m)=\max\{0,S\},$ (5)
where $\eta$ is the acceptance threshold. Multiple runs may be averaged to mitigate variance.
Accepted items are stored together with their initial weights and full provenance (hashes, seeds, versions, and A/B fingerprints), while rejected or ambiguous items are excluded. This procedure yields a compact, auditable admission signal that can be validated offline and efficiently executed in parallel without dependence on the original environment.
3.3 Self-Scheduling in the Memory Controller
The memory controller manages and optimizes the repository through a self-scheduling policy. Unlike traditional systems that depend on costly per-query reranking Nogueira and Cho (2019); Ren et al. (2021); Ni et al. (2022), our approach establishes an evidence-based mechanism for both selecting memory items during retrieval and continuously refining the repository. It comprises two core functions: retrieval-time scheduling, which determines how to use memories effectively for an incoming query, and consolidation and progressive evolution, which curates a compact, high-quality memory set. Together, these components ensure that memory usage is grounded in verifiable utility signals, improving both efficiency and performance.
3.3.1 Retrieval-time Scheduling
The controller’s scheduling policy relies on a ranking signal aligned with realized utility, avoiding the instability and computational cost of on-the-fly large language model reranking Qin et al. (2023). Prior approaches typically use vector similarity or ad-hoc prompt-based scoring, but semantic similarity alone does not guarantee actual task benefit, and repeated reranking adds latency and variance Karpukhin et al. (2020); Izacard and Grave (2021) or ad-hoc prompt-based scoring Nogueira and Cho (2019); Qin et al. (2023).
In our design, we incorporate evidence collected at write time via A/B validation on Self-Contained Execution Contexts (SCECs). These statistics, particularly the measured changes in reward and latency, are mapped into a stable admission-derived weight $w(m)$ for each memory item. At retrieval time, this weight is combined with semantic similarity to form a utility-aligned score:
$s(q,m)=\operatorname{sim}(q,m)\times w(m),$ (6)
where $\operatorname{sim}(q,m)$ denotes the semantic similarity between query $q$ and memory $m$, and $w(m)$ reflects its empirically validated utility.
This coupling of semantic relevance with admission-grounded evidence stabilizes selection and reduces overhead, ensuring that memory items are injected into prompts not only because they are similar, but because they have demonstrated measurable benefit. As a result, retrieval decisions are both efficient and aligned with the system’s long-term objectives.
3.3.2 Consolidation and Progressive Evolution
The consolidation and progressive evolution module maintains a compact yet effective memory repository by suppressing redundancy, preserving items with stable gains, and eliminating or recycling items that show conflicts or sustained negative contributions. While retrieval-time scheduling focuses on selecting memories for specific queries, consolidation and evolution aim to improve the repository itself so that subsequent scheduling operates on a cleaner and more reliable basis.
Progressive evolution is achieved by tracking usage and outcome signals and applying conservative updates to utility weights over time. Items that are rarely retrieved or consistently fail to provide positive utility are gradually decayed, reducing their influence in future selections. Conversely, items that repeatedly yield positive gains across related contexts are promoted. Promotion increases their weights and, when consistency is observed across multiple items, may trigger abstraction into higher-level insights. Such abstractions enable representative entries to replace families of consistent low-level experiences Goyal et al. (2019); Parisotto et al. (2019). If observed outcomes diverge significantly from admission-time evidence, items are demoted or queued for cleanup. All updates are logged with provenance to ensure the evolution process remains auditable Buneman et al. (2001); Lebo et al. (2013); Chirigati et al. (2016).
The weight update for a memory item $m$ depends on its current weight $w(m)$, usage frequency $f_{use}(m)$, and average realized utility $\bar{U}(m)$ since the last update:
$w_{t+1}(m)=w_{t}(m)+\alpha\cdot\bar{U}_{t}(m)-\beta\cdot f_{\mathrm{use},t}(m),$ (7)
where $\alpha$ and $\beta$ control the influence of observed utility and usage. This ensures weights evolve from admission-derived values toward refined, usage-aware estimates.
Conflict detection is integral to this loop. A memory is marked as conflicting when repeated injections consistently reduce task reward or when its implications contradict other rules. Such items undergo progressive weight reduction, and if their weight falls below a threshold, they are demoted or removed. All decisions retain version traces and evidence chains to support rollback and inspection.
Semantic consolidation addresses redundancy among items from different SCECs. When two or more items, $m_{i}$ and $m_{j}$, show high semantic similarity without conflicting applicability, they are merged:
$m_{\text{merged}}=\text{Merge}(m_{i},m_{j}).$ (8)
The merged entry preserves essential content while aggregating evidence from the originals, and its weight $w(m_{\text{merged}})$ is reconciled to reflect combined support without double counting. Contributing items are archived or soft-deleted, preserving provenance if needed. By collapsing near-duplicates into single representatives, the repository reduces retrieval noise and strengthens utility signals Broder et al. (1997); Manku et al. (2007); Charikar (2002); Zhang et al. (2019).
Through these routines, the controller maintains a concise but reliable set of memories. Redundant entries are consolidated Manku et al. (2007), stable positives are promoted Kaelbling et al. (1996), abstractions generalize recurring insights Goyal et al. (2019), and harmful items are isolated or removed Toneva et al. (2019). As a result, the repository remains aligned with realized utility, ensuring that retrieval decisions exploit empirical evidence rather than ad-hoc reranking.
3.4 Cross-Domain Knowledge Diffusion
This component exploits the environment-free and verifiable properties of the SCEC method, treating memory entries as portable and re-verifiable assets across domains. Tasks from diverse domains continuously supply evidence to refine memory weights, thereby improving both universality and robustness of the repository. The process follows a loop of migrate, re-validate, and re-incorporate: transfer is initiated through retrieval and weighted injection; subsequent usage allows re-estimation of weights via SCEC-compatible procedures; and the updates inform later scheduling and admission decisions. Throughout, retrieval signals remain unchanged: similarity-based relevance, $\operatorname{sim}(q,m)$ for query $q$ and memory item $m$, combined with the admission-derived weight $w(m)$. This design avoids evidence-free cold starts and eliminates the need for additional scoring components at runtime.
Immediately after admission, each specific entry $m_{\text{specific}}$ generates a conservative general form $m_{\text{general}}$ through a lightweight abstraction operator:
$m_{\text{general}}=\operatorname{Abstract}(m_{\text{specific}}).$ (9)
This yields a dual-linked pair in which the general form strips domain-specific features while preserving transferable essence. The general form serves as a low-risk candidate for cross-domain retrieval, while the specific form remains primary within its source domain.
The abstraction process is rule-governed and minimal. Entities and domain-specific terms are replaced with typed placeholders, retaining actionable task–action structures while removing non-essential detail Goyal et al. (2019); Lake et al. (2015); Liang et al. (2017). The result is a compact snippet suitable for direct injection and controlled comparisons. To minimize orchestration cost, abstraction is generated alongside the SCEC-based A/B assessment within the distributed pipeline Chirigati et al. (2016).
For weight inheritance, the general form is initialized conservatively as a scaled version of the specific weight, $w_{\text{general}}=\alpha\cdot w_{\text{specific}}$ with $\alpha<1$. This prior encodes caution against over-abstraction while preserving provenance for later auditing and weight updates.
At retrieval, both forms compete in the candidate set with a unified score:
$s(q,m)=\operatorname{sim}(q,m)\times w(m).$ (10)
In-domain queries typically favor specific forms due to higher semantic match and weight, while cross-domain queries benefit from the more stable similarity of general forms. This mechanism enables knowledge diffusion across domains without introducing extra runtime complexity, while maintaining auditability, portability, and stability. Subsequent use further refines the weights, allowing knowledge to propagate adaptively across diverse tasks.
4 Experiment
4.1 Experimental Setup
Datasets.
Experiments are conducted on two widely used benchmark datasets: FEVER(Thorne et al., 2018) and HotpotQA(Yang et al., 2018b).HotpotQA is a large-scale question-answering benchmark designed to evaluate the ability of systems to perform multi-hop reasoning across diverse natural language inputs. FEVER is a fact-checking dataset that provides human-written claims about Wikipedia entities, each labeled as Supported, Refuted, or NotEnoughInfo. Both datasets present significant challenges for testing long-term reasoning and memory utilization in language agents.
Baselines.
The proposed SEDM is compared with the following baselines: (1) No Memory: the model only relies on the query input without any memory augmentation, serving as the basic performance reference; (2) G-Memory(Zhang et al., 2025): a memory-augmented method that stores all past information in a global memory pool and retrieves by similarity search. Although effective, it incurs high inference cost due to the large number of prompt tokens; (3) SEDM (ours): our scalable self-evolving distributed memory, which introduces memory scheduling and selection mechanisms to maintain a compact and adaptive working set, thereby balancing performance and efficiency.
Configurations.
All experiments run on the same backbone LLM (GPT-4o-mini) OpenAI (2023). Dense retrieval is handled by ALL-MINILM-L6-V2 Reimers and Gurevych (2019), which embeds both knowledge snippets and queries for similarity search. On the evaluation side, we use FEVER accuracy for fact-checking and HotpotQA exact-match (EM) for multi-hop QA. Efficiency is tracked by counting prompt and completion tokens consumed during inference. To ensure fair comparisons, every method is granted the same memory budget; the proposed SEDM does not expand this budget, but instead adaptively schedules which entries are kept in memory.
4.2 Result Analysis
The overall performance on FEVER and HotpotQA is summarized in Table 1. In the FEVER dataset, the baseline model achieved only 57 without memory, reflecting limited reasoning ability in the absence of external knowledge and prior memory. G-Memory improved the score to 62, but this gain came at the cost of a dramatic increase in the number of prompt tokens, leading to significantly higher inference costs. In contrast, SEDM achieved the highest score of 66 while consuming far fewer tokens than G-Memory. This demonstrates that our method successfully balances the trade-off between performance and efficiency through its memory selection and scheduling mechanisms.
In the HotpotQA dataset, the trend is similar to that observed in FEVER. The no-memory baseline scored only 34, while G-Memory increased the score to 38. SEDM further improved performance, reaching a score of 39 while simultaneously reducing computational overhead, confirming its effectiveness in multi-hop reasoning tasks.
Moreover, we evaluate the transfer ability of SEDM between FEVER and HotpotQA, two distinct downstream tasks. Specifically, the agent collects experience on the HotpotQA task using SEDM and then evaluates it on FEVER to measure knowledge transfer and prompting effects. Under this setting, the score on FEVER reached 64. Compared with G-Memory, which scored 62, and the no-memory baseline, which scored 57, our results demonstrate that SEDM enables adaptive memory selection that leverages previously collected experiences to improve performance across tasks.
Table 1: Performance comparison on FEVER (fact verification) and HotpotQA (multi-hop reasoning). We report task accuracy (Score) along with efficiency metrics (Prompt Tokens and Completion Tokens). SEDM achieves the best accuracy on both benchmarks while substantially reducing token consumption compared with G-Memory, highlighting its ability to balance effectiveness and efficiency.
Method FEVER HotpotQA
Score Prompt Tokens Completion Tokens Score Prompt Tokens Completion Tokens
No Memory 57 1.65M 24K 34 2.46M 29K
G-Memory 62 3.62M 109K 38 4.63M 114K
SEDM (Ours) 66 2.47M 53K 39 3.88M 55K
Table 2: Ablation study on HotpotQA and FEVER, showing the progressive contribution of SEDM components.
Dataset Setting Score Prompt tokens Completion tokens
HotpotQA No Memory 34 2.46M 29K
+ SCEC 37 3.52M 52K
+ SCEC + Self-Scheduling 39 3.88M 55K
FEVER No Memory 57 1.65M 24K
+ SCEC 64 2.19M 53K
+ SCEC + Self-Scheduling 66 2.47M 53K
4.3 Ablation Study
To evaluate the contribution of individual SEDM components, we conduct ablation studies on both HotpotQA and FEVER. Table 2 reports results under three configurations: (i) the baseline without memory, (ii) the addition of SCEC-based verifiable write admission (+SCEC), and (iii) the full SEDM with the memory controller’s self-scheduling mechanism (+SCEC + Self-Scheduling).
On HotpotQA, introducing +SCEC improves the score from 34 to 37, but also increases prompt tokens by 43% (2.46M $\rightarrow$ 3.52M) and completion tokens from 29K to 52K. With +Self-Scheduling, the score further rises to 39, while prompt tokens grow only by 10% (3.52M $\rightarrow$ 3.88M), showing that scheduling effectively controls token overhead relative to the accuracy gain. On FEVER, the baseline achieves 57. Adding +SCEC raises the score to 64, accompanied by an increase in prompt tokens from 1.65M to 2.19M (+33%) and completion tokens from 24K to 53K. With scheduling, performance improves to 66, while prompt tokens rise only to 2.47M (+13%), and completion tokens remain unchanged, confirming that the controller filters relevant memory without inflating responses.
In summary, across both datasets, SCEC consistently yields substantial accuracy gains at the cost of increased token usage, while the self-scheduling mechanism provides further improvements with relatively minor overhead. This demonstrates that SEDM not only enhances reasoning accuracy but also achieves a more favorable trade-off between performance and efficiency.
4.4 Cross-Domain Evaluation
To further assess the generalization ability of SEDM across domains, we conduct a cross-domain experiment in which memory is collected on one dataset and evaluated on another. Table 3 reports the results on both FEVER and HotpotQA.
Table 3: Cross-domain evaluation of SEDM. Rows indicate the dataset used for memory collection, and columns indicate the dataset used for testing.
Collect $\downarrow$ / Test $\rightarrow$ Fever HotpotQA
Fever 66 41
HotpotQA 64 39
The results highlight two key findings. First, in-domain settings (diagonal entries) achieve the strongest performance, with scores of 66 on FEVER and 39 on HotpotQA, confirming that SEDM effectively distills and leverages domain-specific memory. Second, cross-domain transfer (off-diagonal entries) reveals an interesting asymmetry. When memory distilled from FEVER is applied to HotpotQA, the model achieves a score of 41, which even surpasses the in-domain HotpotQA result (39). This indicates that factual knowledge captured in FEVER has strong generalization capacity and can support reasoning tasks that require grounding in factual claims. By contrast, memory distilled from HotpotQA transfers less effectively to FEVER (64 vs. 66), suggesting that multi-hop reasoning knowledge is less directly reusable for fact verification. Overall, these findings demonstrate that SEDM not only excels in in-domain scenarios but also exhibits promising cross-domain generalization, particularly when transferring from simpler fact verification tasks to more complex multi-hop reasoning tasks.
5 Conclusion
This paper introduces SEDM, Scalable Self-Evolving Distributed Memory, which transforms memory in multi-agent systems from a passive repository into an adaptive and verifiable component by integrating SCEC-based admission, self-scheduling refinement, and cross-domain knowledge diffusion. Through this principled design, SEDM addresses the challenges of noise accumulation, uncontrolled growth, and weak generalization that limit existing methods. Experiments on FEVER and HotpotQA confirm that SEDM improves reasoning accuracy while reducing computational and token overhead, demonstrating its potential as a scalable and sustainable memory mechanism for long-term multi-agent collaboration.

### v3

@@
 Abstract
-Long-term multi-agent systems inevitably generate vast amounts of trajectories and historical interactions, which makes efficient memory management essential for both performance and scalability.
-Existing methods typically depend on vector retrieval and hierarchical storage, yet they are prone to noise accumulation, uncontrolled memory expansion, and limited generalization across domains.
-To address these challenges, we present SEDM, Self-Evolving Distributed Memory, a verifiable and adaptive framework that transforms memory from a passive repository into an active, self-optimizing component.
+In long-term multi-agent systems, the accumulation of trajectories and historical interactions makes efficient memory management a particularly challenging task, with significant implications for both performance and scalability.
+Existing memory management methods typically depend on vector retrieval and hierarchical storage, yet they are prone to noise accumulation, uncontrolled memory expansion, and limited generalization across domains.
+To address these challenges, we present SEDM (Self-Evolving Distributed Memory), a verifiable and adaptive framework that transforms memory from a passive repository into an active, self-optimizing component.
 SEDM integrates verifiable write admission based on reproducible replay, a self-scheduling memory controller that dynamically ranks and consolidates entries according to empirical utility, and cross-domain knowledge diffusion that abstracts reusable insights to support transfer across heterogeneous tasks.
 Evaluations on benchmark datasets demonstrate that SEDM improves reasoning accuracy while reducing token overhead compared with strong memory baselines, and further enables knowledge distilled from fact verification to enhance multi-hop reasoning.
-The results highlight SEDM as a scalable and sustainable memory mechanism for open-ended multi-agent collaboration.
-The code will be released in the later stage of this project.
+The results highlight SEDM as a scalable and sustainable memory management design for open-ended multi-agent collaboration.
+The code will be released upon acceptance of this paper.
 1 Introduction
-In recent years, the rapid development of large-scale multi-agent systems (MAS) Wooldridge (2009); Lowe et al.
-(2017); Foerster et al.
-(2016); Yang et al.
-(2018a); Buşoniu et al.
-(2008).
-has expanded their application in diverse domains, including collaborative reasoning, decision-making, and autonomous planning Shoham et al.
-(2007).
-A central challenge in these open-ended and long-term tasks is the ability of agents to manage, interpret, and reuse information accumulated from continuous interaction with both peers and the environment Yang et al.
-(2018a).
-Without effective memory mechanisms, the sheer scale of historical interactions can easily overwhelm computational resources and compromise decision quality Foerster et al.
-(2018).
-In open-ended and long-term multi-agent tasks, each agent relies on its past memories, the observed states of other agents, and the current environment to determine subsequent actions or responses Das et al.
-(2017).
-During continuous interaction between agents and their environment, the MAS gradually accumulates extensive logs of interactions, invocation trajectories, and high-level policy memories OpenAI (2023).
-Such overwhelming amounts of information directly impact the efficiency and cost of decision-making, often leading to higher monetary costs and longer contextual requirements for inference Shuster et al.
-(2020).
+In recent years, the rapid development of large-scale multi-agent systems (MAS) Wooldridge2009MAS ; Lowe2017MADDPG ; Foerster2016Comm ; cite2 ; Busoniu2008Survey has expanded their application in diverse domains, including collaborative reasoning, decision-making, and autonomous planning cite1 .
+A fundamental challenge in open-ended, long-term tasks lies in enabling agents to effectively manage, interpret, and reutilize information accumulated through continuous interactions with both peers and their environment cite2 .
+In the absence of effective memory management design, the sheer scale of historical interactions can easily overwhelm computational resources and compromise decision quality cite3 .
+In open-ended and long-term multi-agent tasks, each agent relies on its past memories, the observed states of other agents, and the current environment to make decisions for subsequent actions or responses cite4 .
+During continuous interaction between agents and their environment, the MAS gradually accumulates extensive logs of interactions, invocation trajectories, and high-level policy memories openai2024gpt4 .
+Such overwhelming amounts of information directly impact the efficiency and cost of decision-making, often leading to higher monetary costs and longer contextual requirements for inference cite6 .
 Therefore, designing an efficient and sustainable memory mechanism has become a critical issue for modern long-term multi-agent systems.
-Current methods primarily adopt vector retrieval and hierarchical memory structures to manage storage and retrieval efficiently Johnson et al.
-(2019b).
-Vector retrieval Johnson et al.
-(2019a); Karpukhin et al.
-(2020); Izacard and Grave (2021); Lewis et al.
-(2020); Guo et al.
-(2016) leverages semantic similarity to identify relevant entries, while hierarchical organization arranges information in layered structures according to abstraction levels Rae et al.
-(2021).
-These approaches have shown promise in improving retrieval accuracy and managing memory scalability Chen et al.
-(2016).
-However, in complex collaborative multi-agent tasks, their effectiveness diminishes, as the underlying assumptions of stability and linear growth do not hold Wang et al.
-(2020).
+Current methods primarily adopt vector retrieval and hierarchical memory structures to manage storage and retrieval efficiently cite8 .
+Vector retrieval Johnson2019FAISS ; cite34 ; cite33 ; cite14 ; Guo2016DRMM leverages semantic similarity to identify relevant entries, while hierarchical organization arranges information in layered structures according to abstraction levels cite9 .
+These approaches have shown promise in improving retrieval accuracy and managing memory scalability cite10 .
+However, in complex collaborative multi-agent tasks, their effectiveness diminishes, as the underlying assumptions of stability and linear growth do not hold cite11 .
 This gap between theoretical promise and practical performance highlights several critical limitations that hinder their long-term applicability.
-One major challenge is the inevitable accumulation of noise, which severely degrades retrieval quality Fan et al.
-(2018).
-As the memory size expands without constraint, the system faces exponentially increasing computational costs in both retrieval and context construction Izacard and Grave (2021).
-This not only reduces overall efficiency but also amplifies the interference caused by redundant information Lewis et al.
-(2020).
-In particular, the presence of low-value or semantically irrelevant entries dilutes the contribution of high-quality information in retrieval results, impairing downstream task performance and leading to measurable declines in metrics Chen et al.
-(2017).
-In addition, the cumulative noise effect increases response latency and accelerates the nonlinear consumption of computational and storage resources Karpukhin et al.
-(2020), ultimately threatening both scalability and stability in long-term MAS operations Press et al.
-(2020).
+One major challenge is the inevitable accumulation of noise, which severely degrades retrieval quality cite12 .
+As the memory size expands without constraint, the system faces exponentially increasing computational costs in both retrieval and context construction cite33 .
+This not only reduces overall efficiency but also amplifies the interference caused by redundant information cite14 .
+In particular, the presence of low-value or semantically irrelevant entries dilutes the contribution of high-quality information in retrieval results, impairing downstream task performance and leading to measurable declines in metrics cite15 .
+In addition, the cumulative noise effect increases response latency and accelerates the nonlinear consumption of computational and storage resources cite34 , ultimately threatening both scalability and stability in long-term MAS operations cite17 .
 To overcome these limitations, we introduce Scalable Self-Evolving Distributed Memory (SEDM), a framework that transforms memory from a passive repository into an adaptive, self-optimizing, and verifiable component for multi-agent systems.
@@
 It operationalizes memory as an active mechanism by integrating verifiability and continuous self-improvement into the memory lifecycle.
-At its core, memory items undergo a rigorous admission process based on self-contained execution contexts (SCECs), such as Docker and ReproZip Merkel (2014); Chirigati et al.
-(2016), which package all necessary information for environment-free replay and offline validation.
+At its core, memory items undergo a rigorous admission process based on self-contained execution contexts (SCECs), such as Docker and ReproZip Merkel2014Docker ; chirigati2016reprozip , which package all necessary information for environment-free replay and offline validation.
 This mechanism provides empirical evidence for utility at write time, ensuring that only useful, high-quality experiences enter the memory repository.
@@
 Together, these components establish a scalable and auditable memory mechanism that enhances reasoning accuracy, reduces overhead, and supports sustainable long-term multi-agent collaboration.
-We evaluate SEDM on two representative benchmarks, FEVER Thorne et al.
-(2018) for fact verification and HotpotQA Yang et al.
-(2018b) for multi-hop reasoning, comparing against no-memory and G-Memory baselines Zhang et al.
-(2025).
+We evaluate SEDM on two representative benchmarks, FEVER thorne-etal-2018-fever for fact verification and HotpotQA yang2018hotpotqadatasetdiverseexplainable for multi-hop reasoning, comparing against no-memory and G-Memory baselines zhang2025gmemorytracinghierarchicalmemory .
 The results show that SEDM consistently improves task accuracy while significantly reducing token overhead, thereby achieving a better balance between performance and efficiency.
@@
 Our contributions are summarized as follows:
-  • We propose Scalable Self-Evolving Distributed Memory (SEDM), a novel framework that transforms memory from a passive repository into an adaptive, verifiable, and continuously improving component, introducing self-contained execution contexts (SCECs) for reproducible admission and utility-based memory weighting.
+  • We propose Self-Evolving Distributed Memory (SEDM), a novel framework that transforms memory from a passive repository into an adaptive, verifiable, and continuously improving component, introducing self-contained execution contexts (SCECs) for reproducible admission and utility-based memory weighting.
   • We design a self-scheduling memory controller that selectively manages memory at retrieval time and continuously refines the repository through consolidation, redundancy suppression, and progressive evolution, thereby balancing accuracy and efficiency.
-  • We conduct extensive evaluations on FEVER and HotpotQA benchmarks, demonstrating that SEDM consistently improves task accuracy while significantly reducing token overhead.
+  • We conduct extensive evaluations on LoCoMo, FEVER, and HotpotQA benchmarks, demonstrating that SEDM consistently improves task accuracy while significantly reducing token overhead.
 2 Related Work
@@
 Recent efforts in building self-evolving agents have focused on enabling systems to improve their reasoning or behavior over time without explicit retraining.
-Approaches such as Reflexion Shinn et al.
-(2023) and Voyager Wang et al.
-(2023) allow agents to iteratively refine their strategies by leveraging self-reflection and accumulated trajectories.
-Similarly, MEMIT Meng et al.
-(2022) demonstrates the feasibility of localized knowledge editing within large language models, suggesting a pathway for agents to evolve by continuously updating their internal representations.
+Approaches such as Reflexion shinn2023reflexion and Voyager wang2023voyager allow agents to iteratively refine their strategies by leveraging self-reflection and accumulated trajectories.
+Similarly, MEMIT meng2022memit demonstrates the feasibility of localized knowledge editing within large language models, suggesting a pathway for agents to evolve by continuously updating their internal representations.
 These studies highlight the importance of mechanisms that support autonomous adaptation and progressive self-improvement in dynamic environments.
@@
 Episodic memory systems, such as those proposed by Park et al.
-Park et al.
-(2023), emulate human-like memory consolidation to support consistent long-term behavior in simulated social environments.
-Memory-augmented neural networks Graves et al.
-(2016) and differentiable neural dictionaries Kaiser et al.
-(2017) further demonstrate how structured memory access can enhance reasoning and generalization.
-More recently, retrieval-augmented generation frameworks tailored for interactive agents Mialon et al.
-(2023) have shown that dynamically grounding responses in curated external memories improves both interpretability and task success.
+park2023generativeagents , emulate human-like memory consolidation to support consistent long-term behavior in simulated social environments.
+Memory-augmented neural networks graves2016hybrid and differentiable neural dictionaries kaiser2017learning further demonstrate how structured memory access can enhance reasoning and generalization.
+More recently, retrieval-augmented generation frameworks tailored for interactive agents mialon2023augmented have shown that dynamically grounding responses in curated external memories improves both interpretability and task success.
 Together, these works underscore the need for memory systems that are not only scalable but also adaptive to the agent’s evolving operational context.
@@
 The memory controller manages and optimizes the repository through a self-scheduling policy.
-Unlike traditional systems that depend on costly per-query reranking Nogueira and Cho (2019); Ren et al.
-(2021); Ni et al.
-(2022), our approach establishes an evidence-based mechanism for both selecting memory items during retrieval and continuously refining the repository.
+Unlike traditional systems that depend on costly per-query reranking nogueira2019passage ; ren2021rocketqav2 ; ni2022large , our approach establishes an evidence-based mechanism for both selecting memory items during retrieval and continuously refining the repository.
 It comprises two core functions: retrieval-time scheduling, which determines how to use memories effectively for an incoming query, and consolidation and progressive evolution, which curates a compact, high-quality memory set.
@@
 3.3.1 Retrieval-time Scheduling
-The controller’s scheduling policy relies on a ranking signal aligned with realized utility, avoiding the instability and computational cost of on-the-fly large language model reranking Qin et al.
-(2023).
-Prior approaches typically use vector similarity or ad-hoc prompt-based scoring, but semantic similarity alone does not guarantee actual task benefit, and repeated reranking adds latency and variance Karpukhin et al.
-(2020); Izacard and Grave (2021) or ad-hoc prompt-based scoring Nogueira and Cho (2019); Qin et al.
-(2023).
+The controller’s scheduling policy relies on a ranking signal aligned with realized utility, avoiding the instability and computational cost of on-the-fly large language model reranking wu2023llmrerank .
+Prior approaches typically use vector similarity or ad-hoc prompt-based scoring, but semantic similarity alone does not guarantee actual task benefit, and repeated reranking adds latency and variance cite34 ; cite33 or ad-hoc prompt-based scoring nogueira2019passage ; wu2023llmrerank .
 In our design, we incorporate evidence collected at write time via A/B validation on Self-Contained Execution Contexts (SCECs).
@@
 Promotion increases their weights and, when consistency is observed across multiple items, may trigger abstraction into higher-level insights.
-Such abstractions enable representative entries to replace families of consistent low-level experiences Goyal et al.
-(2019); Parisotto et al.
-(2019).
+Such abstractions enable representative entries to replace families of consistent low-level experiences goyal2019abstraction ; parisotto2019stabilizing .
 If observed outcomes diverge significantly from admission-time evidence, items are demoted or queued for cleanup.
-All updates are logged with provenance to ensure the evolution process remains auditable Buneman et al.
-(2001); Lebo et al.
-(2013); Chirigati et al.
-(2016).
+All updates are logged with provenance to ensure the evolution process remains auditable buneman2001provenance ; groth2012provenance ; chirigati2016reprozip .
 The weight update for a memory item $m$ depends on its current weight $w(m)$, usage frequency $f_{use}(m)$, and average realized utility $\bar{U}(m)$ since the last update:
@@
 Contributing items are archived or soft-deleted, preserving provenance if needed.
-By collapsing near-duplicates into single representatives, the repository reduces retrieval noise and strengthens utility signals Broder et al.
-(1997); Manku et al.
-(2007); Charikar (2002); Zhang et al.
-(2019).
+By collapsing near-duplicates into single representatives, the repository reduces retrieval noise and strengthens utility signals broder1997resemblance ; manku2007detecting ; charikar2002simhash ; zhang2019duplicate .
 Through these routines, the controller maintains a concise but reliable set of memories.
-Redundant entries are consolidated Manku et al.
-(2007), stable positives are promoted Kaelbling et al.
-(1996), abstractions generalize recurring insights Goyal et al.
-(2019), and harmful items are isolated or removed Toneva et al.
-(2019).
+Redundant entries are consolidated manku2007detecting , stable positives are promoted kaelbling1996rlsurvey , abstractions generalize recurring insights goyal2019abstraction , and harmful items are isolated or removed toneva2019catastrophic .
 As a result, the repository remains aligned with realized utility, ensuring that retrieval decisions exploit empirical evidence rather than ad-hoc reranking.
@@
 The abstraction process is rule-governed and minimal.
-Entities and domain-specific terms are replaced with typed placeholders, retaining actionable task–action structures while removing non-essential detail Goyal et al.
-(2019); Lake et al.
-(2015); Liang et al.
-(2017).
+Entities and domain-specific terms are replaced with typed placeholders, retaining actionable task–action structures while removing non-essential detail goyal2019abstraction ; lake2015human ; dong2019neural .
 The result is a compact snippet suitable for direct injection and controlled comparisons.
-To minimize orchestration cost, abstraction is generated alongside the SCEC-based A/B assessment within the distributed pipeline Chirigati et al.
-(2016).
+To minimize orchestration cost, abstraction is generated alongside the SCEC-based A/B assessment within the distributed pipeline chirigati2016reprozip .
 For weight inheritance, the general form is initialized conservatively as a scaled version of the specific weight, $w_{\text{general}}=\alpha\cdot w_{\text{specific}}$ with $\alpha<1$.
@@
 4.1 Experimental Setup
-Datasets.
-Experiments are conducted on two widely used benchmark datasets: FEVER(Thorne et al., 2018) and HotpotQA(Yang et al., 2018b).HotpotQA is a large-scale question-answering benchmark designed to evaluate the ability of systems to perform multi-hop reasoning across diverse natural language inputs.
+4.1.1 Dataset and Model
+Evaluation is conducted on the LoCoMo benchmark maharana2024LoCoMo , which consists of two components: (i) multi-turn dialogues (approximately 600 turns per dialogue, averaging 26,000 tokens) and (ii) question-answer (QA) pairs grounded in those dialogues.
+Each dialogue contains roughly 200 questions spanning single-hop rajpurkar2016squad , multi-hop yang2018hotpotqa , open-domain kwiatkowski2019natural , and temporal reasoning chen2021timeqa .
+All experiments are carried out with gpt-4o-mini to ensure consistency and comparability across evaluations.
+Performance is measured using two complementary metrics: Token-level F1 Score (F1) rajpurkar2016squad , which captures the overlap between predicted and ground-truth answers, and BLEU-1 (B1) papineni2002bleu , which evaluates unigram-level lexical similarity.
+4.1.2 Baselines
+To evaluate the effectiveness of our framework SEDM, we compare it with several representative baselines for multi-session dialogue reasoning: LoCoMo maharana2024LoCoMo , a benchmark for assessing long-range retrieval and reasoning in multi-session conversations; ReadAgent lee2024human , a human-in-the-loop agent for long-context reading; MemoryBank zhong2024memorybank , a memory-augmented retrieval model; MemoryGPT packer2024memgptllmsoperatingsystems , an LLM operating system with hierarchical memory modules; A-Mem xu2025amemagenticmemoryllm , a dynamic agentic memory system that creates, links, and updates structured memories; Zep rasmussen2025zeptemporalknowledgegraph , a retrieval-based agent with structured memory access for temporally extended queries; LangMem wang2023augmentinglanguagemodelslongterm , an open-source framework connecting memory chains across sessions; Mem0 chhikara2025mem0buildingproductionreadyai , a modular memory system with explicit in-context memory operations; and G-memory zhang2025gmemorytracinghierarchicalmemory , which leverages hierarchical tracing for memory retrieval.
+4.2 Main results
+Table 1: Evaluation results on the LoCoMo benchmark dataset comparing SEDM with other memory-enabled systems.
+Models are evaluated on F1 rajpurkar2016squad and BLEU-1 (B1) papineni2002bleu across Single Hop rajpurkar2016squad , Multi-Hop yang2018hotpotqa , Open Domain kwiatkowski2019natural , Temporal, and Adversial questions chen2021timeqa .
+Higher is better.
+The top two best results are marked in Bold.
+Method Single Hop Multi-Hop Open Domain Temporal Adversial
+[…]
+LoCoMo maharana2024LoCoMo 25.0 19.8 18.4 14.8 40.4 29.1 12.0 11.2 69.2 68.8
+ReadAgent lee2024human 9.2 6.5 12.6 8.9 9.7 7.7 5.3 5.1 9.8 9.0
+MemoryBank zhong2024memorybank 5.0 4.8 9.7 7.0 6.6 5.2 5.6 5.9 7.4 6.5
+MemoryGPT packer2024memgptllmsoperatingsystems 26.7 17.7 25.5 19.4 41.0 34.3 9.2 7.4 43.2 42.7
+A-Mem xu2025amemagenticmemoryllm 27.0 20.1 45.9 36.7 44.7 37.1 12.1 12.0 50.0 49.5
+Zep rasmussen2025zeptemporalknowledgegraph 30.2 17.2 15.0 11.6 26.7 18.4 3.5 2.7 22.6 15.1
+LangMem langchain2024 22.4 15.2 18.7 16.0 31.6 23.9 27.8 21.5 28.3 21.3
+Mem0 chhikara2025mem0buildingproductionreadyai 27.3 18.6 18.6 13.9 34.0 24.8 26.9 21.1 30.4 22.2
+G-memory zhang2025gmemorytracinghierarchicalmemory 34.6 26.6 9.05 7.2 53.5 44.0 32.4 25.6 11.3 9.3
+SEDM (Ours) 33.5 24.4 12.1 9.2 51.7 37.0 47.5 33.1 12.1 9.3
+Table 1 presents the results of SEDM compared with strong baselines, including LoCoMo, ReadAgent, MemoryBank, MemoryGPT, A-Mem, Zep, LangMem, Mem0, and G-memory, on the LoCoMo benchmark across Single Hop, Multi-Hop, Open Domain, Temporal, and Adversarial reasoning tasks.
+Overall, the results show that SEDM delivers strong and consistent improvements in challenging reasoning scenarios.
+In particular, SEDM achieves the highest F1 and BLEU-1 scores on the Open Domain and Temporal settings, surpassing all other baselines by a substantial margin.
+For instance, SEDM improves Temporal reasoning by more than $15$ F1 points compared with the strongest baseline (G-memory), demonstrating its ability to capture and utilize temporally structured information effectively.
+On the Single Hop setting, SEDM remains highly competitive, ranking among the top two models alongside G-memory, while showing significantly better balance across different question types.
+Although A-Mem and MemoryGPT achieve strong results on Multi-Hop and Adversarial tasks respectively, their performance drops markedly on other reasoning categories, whereas SEDM maintains stable and robust performance across the board.
+These results highlight that the learned selective memory mechanism in SEDM provides clear advantages over heuristic, retrieval-based, and hierarchical memory systems.
+By selectively integrating context across sessions, SEDM enables LLMs to reason more effectively over long, multi-session dialogues and demonstrates strong generalization across diverse reasoning categories.
+4.3 Efficient analysis on Fever and HotpotQA
+We further conduct Experiments on two widely used benchmark datasets: FEVER(thorne-etal-2018-fever, ) and HotpotQA(yang2018hotpotqadatasetdiverseexplainable, ).HotpotQA is a large-scale question-answering benchmark designed to evaluate the ability of systems to perform multi-hop reasoning across diverse natural language inputs.
 FEVER is a fact-checking dataset that provides human-written claims about Wikipedia entities, each labeled as Supported, Refuted, or NotEnoughInfo.
 Both datasets present significant challenges for testing long-term reasoning and memory utilization in language agents.
-Baselines.
-The proposed SEDM is compared with the following baselines: (1) No Memory: the model only relies on the query input without any memory augmentation, serving as the basic performance reference; (2) G-Memory(Zhang et al., 2025): a memory-augmented method that stores all past information in a global memory pool and retrieves by similarity search.
+The proposed SEDM is compared with the following baselines: (1) No Memory: the model only relies on the query input without any memory augmentation, serving as the basic performance reference; (2) G-Memory(zhang2025gmemorytracinghierarchicalmemory, ): a memory-augmented method that stores all past information in a global memory pool and retrieves by similarity search.
 Although effective, it incurs high inference cost due to the large number of prompt tokens; (3) SEDM (ours): our scalable self-evolving distributed memory, which introduces memory scheduling and selection mechanisms to maintain a compact and adaptive working set, thereby balancing performance and efficiency.
-Configurations.
-All experiments run on the same backbone LLM (GPT-4o-mini) OpenAI (2023).
-Dense retrieval is handled by ALL-MINILM-L6-V2 Reimers and Gurevych (2019), which embeds both knowledge snippets and queries for similarity search.
+All experiments run on the same backbone LLM (GPT-4o-mini) openai2024gpt4 .
+Dense retrieval is handled by ALL-MINILM-L6-V2 reimers2019sentencebert , which embeds both knowledge snippets and queries for similarity search.
 On the evaluation side, we use FEVER accuracy for fact-checking and HotpotQA exact-match (EM) for multi-hop QA.
@@
 To ensure fair comparisons, every method is granted the same memory budget; the proposed SEDM does not expand this budget, but instead adaptively schedules which entries are kept in memory.
-4.2 Result Analysis
 The overall performance on FEVER and HotpotQA is summarized in Table 2.
@@
 To further assess the generalization ability of SEDM across domains, we conduct a cross-domain experiment in which memory is collected on one dataset and evaluated on another.
-Table 3 reports the results on both FEVER and HotpotQA.
+Table 4 reports the results on FEVER, HotpotQA, and LoCoMo.
 Table 4: Cross-domain evaluation of SEDM. Rows indicate the dataset used for memory collection, and columns indicate the dataset used for testing.
-Collect $\downarrow$ / Test $\rightarrow$ Fever HotpotQA
-Fever 66 41
-HotpotQA 64 39
-The results highlight two key findings.
-First, in-domain settings (diagonal entries) achieve the strongest performance, with scores of 66 on FEVER and 39 on HotpotQA, confirming that SEDM effectively distills and leverages domain-specific memory.
-Second, cross-domain transfer (off-diagonal entries) reveals an interesting asymmetry.
-When memory distilled from FEVER is applied to HotpotQA, the model achieves a score of 41, which even surpasses the in-domain HotpotQA result (39).
-This indicates that factual knowledge captured in FEVER has strong generalization capacity and can support reasoning tasks that require grounding in factual claims.
-By contrast, memory distilled from HotpotQA transfers less effectively to FEVER (64 vs.
-66), suggesting that multi-hop reasoning knowledge is less directly reusable for fact verification.
-Overall, these findings demonstrate that SEDM not only excels in in-domain scenarios but also exhibits promising cross-domain generalization, particularly when transferring from simpler fact verification tasks to more complex multi-hop reasoning tasks.
+Collect $\downarrow$ / Test $\rightarrow$ FEVER HotpotQA LoCoMo
+FEVER 66 41 38.1
+HotpotQA 64 39 38.6
+LoCoMo 65 34 37.6
+FEVER as target: The best score is achieved by FEVER itself (66); memory from HotpotQA and LoCoMo yields slightly lower scores (64 and 65), indicating that fact-verification benefits most from in-domain memory.
+HotpotQA as target: The highest score is obtained by FEVER → HotpotQA (41), 2 points above the in-domain result (39).
+Memory from LoCoMo drops to 34, the lowest cross-domain score, suggesting that dialogue-grounded knowledge is least useful for multi-hop reasoning.
+LoCoMo as target: All three sources perform within 1 point of each other (37.6–38.6). Thus, no single source dominates, and dialogue-grounded evaluation is remarkably robust to the origin of memory.
+Overall, SEDM exhibits task-dependent transfer: factual-verification memory transfers surprisingly well to HotpotQA, whereas dialogue memory transfers poorly to the other two tasks.
+In-domain memory is not universally optimal, and the benefit of domain alignment varies significantly by task pair.
 5 Conclusion
@@
 Through this principled design, SEDM addresses the challenges of noise accumulation, uncontrolled growth, and weak generalization that limit existing methods.
-Experiments on FEVER and HotpotQA confirm that SEDM improves reasoning accuracy while reducing computational and token overhead, demonstrating its potential as a scalable and sustainable memory mechanism for long-term multi-agent collaboration.
+Experiments on LoCoMo, FEVER, and HotpotQA confirm that SEDM improves reasoning accuracy while reducing computational and token overhead, demonstrating its potential as a scalable and sustainable memory mechanism for long-term multi-agent collaboration.